A car radiator is filled with 4 gal of a 30% antifreeze solution. How much pure antifreeze must be added to have a 75% antifreeze solution?
Answer by stanbon(75887)

A car radiator is filled with 4 gal of a 30% antifreeze solution. How much pure antifreeze must be added to have a 75% antifreeze solution?
-----------------------
Equation:
anti + anti = anti
0.30*4 + 1.00*x = 0.75(4+x)
----
Multiply thru by 100 to get:
-----------------
30*4 + 100x = 75*4 + 75x
--------
25x = 45*4
x = (9/5)4
x = 36/5
---
x = 7 1/5 gallons (amt. of pure antifreeze to add)
----------------
Cheers,
Stan H.
============